# Analyzing Classroom Interaction Data Using Prompt Engineering and Network Analysis

### v1

@@ 2.3 Utterance Classification Results
2.4 Validation of Utterance Classification
For the validation of our prompt engineering approach, we focused on assessing the alignment between intended classifications and actual outcomes. Due to the absence of pre-labeled data in our classroom dialogue dataset, we utilized the DailyDialog dataset (Li et al., 2017) as our external validation source. This dataset provides a structured framework for dialogue act classification, categorizing utterances into four distinct types: inform, questions, directives, and commissive. To evaluate our methodology, we developed dialogue act classification (DAC) specific prompts for this dataset, incorporating five dialogue chunks examples to create few-shot examples prompts. We compared the performance of these few-shot prompts against zero-shot approaches, enabling us to quantify the impact of including dialogue examples in the classification process.
To establish the effectiveness and generalizability of our prompt engineering approach, we conducted comprehensive comparisons with established deep learning models. Context-Aware Self-Attention model (Raheja and Tetreault, 2019, CASA) employs a self-attention mechanism integrated with hierarchical deep neural networks to recognize dialogue acts within conversational contexts. The Deep Co-Interactive Relation Network (Qin et al., 2020, DCR-Net) employs a hierarchical encoder and relation layer to integrate contextual and interaction information for dialogue act classification. A co-interactive graph attention network model (Qin et al., 2021, Co-GAT) jointly performs DAC and sentiment classification, leveraging both contextual and interaction information through cross-utterance and cross-task connections. Additionally, Zhao et al. (2023) presented classification results using prompt engineering with the GPT-3.5-turbo model, and we compared our results against these presented results. We evaluated classification performance using three complementary metrics: macro-F1, weighted-F1 scores, and overall accuracy. Macro-F1 computes the F1 score for each class individually and then averages them, providing a balanced measure of performance across all classes, regardless of their frequencies. This is particularly useful in imbalanced datasets, where certain classes may dominate. On the other hand, Weighted-F1 calculates the F1 score for each class and takes into account the number of true instances for each class, giving more importance to classes with more instances. These metrics help assess the model’s performance from different perspectives, considering both class balance and overall accuracy.
  Model Acc Macro-F1 Weighted-F1
  CASA - 0.78 -
DCR-Net + Co-Attention - 0.79 -
Co-GAT - 0.79 -
ChatGPT 1-shot + P.E 0.67 - 0.65
ChatGPT 3-shot + P.E 0.73 - 0.71
Llama-3.1-8B-Instruct Zero-shot 0.69 0.57 0.68
Llama-3.1-8B-Instruct Few-shot 0.75 0.66 0.74
Table 5: Comparison of Prompt Engineering results with other baselines on the DAC task. Experimental results for CASA, DCR-Net + Co-Attention, and Co-GAT are partly cited from Qin et al. (2021) and Zhao et al. (2023). P.E. stands for prompt engineering where the 1-shot and 3-shot settings use 1 and 3 examples in the prompt, respectively.
Table 5 presents our comparative analysis of prompt engineering approaches and state-of-the-art deep learning models. Results demonstrate that the few-shot prompt engineering method, incorporating five dialogue chunks as examples, achieves substantially improved performance compared to the zero-shot approach. While some comparative studies did not report all evaluation metrics, we have calculated a comprehensive set of metrics for our prompt engineering results to facilitate thorough comparison across different models. This analytical approach ensures equitable performance assessment across varying methodologies.
The comparative analysis reveals that prompt engineering performance, while effective, may not fully match that of deep learning models due to fundamental methodological distinctions. Deep learning approaches benefit from extensive optimization using pre-constructed labeled datasets and generate specialized embeddings tailored to specific tasks. In contrast, prompt engineering operates solely through label descriptions without prior training data, relying on the inherent language understanding capabilities of the model. This distinction becomes particularly significant in contexts where classification criteria exhibit variability due to multiple annotators or crowdsourcing processes. Nevertheless, these differences suggest that prompt engineering may be especially valuable in generating consistent labels, as it helps maintain classification criteria across datasets.
@@ 3.1 Network Construction from Classified Utterances
We formalized student interactions using an adjacency matrix, where the matrix elements indicate the frequency of specific interaction types between two students during the observation period. We then visualized the constructed network structure by using nodes to represent individual students and directed weighted edges between nodes to represent the nature and frequency of the interactions between the two students. For instance, when a student offers multiple explanations to a peer, a directed edge is formed from the explaining student (source) to the listening student (target), with the edge weight reflecting the total number of explanatory interactions. Engagement interactions similarly generate directed edges from the engaging student to the idea originator, with edge weights reflecting interaction frequency. To capture the varying intensity of student engagement more precisely, we implemented a differential weighting scheme for these interactions. While explanatory interactions maintained equal weights, we applied multiplication factors to different levels of engagement: medium-level engagements received double weight, and high-level engagements received triple weight relative to low-level engagements. This weighted approach ensures that more substantive forms of engagement exert proportionally greater influence in the network structure, better reflecting their educational significance.
@@ 3.1.1 Network Graphs
Figure 2 illustrates these network transformations, where each node represents a student and edges represent their interactions. Node colors indicate students’ standardized mathematics performance, utilizing a spectrum from blue (lower scores) to red (higher scores). Edge thickness corresponds to interaction frequency between student pairs.
@@
We then examined network centrality, using metrics designed for directed weighted networks, accounting for both directional flow and interaction intensity. These include PageRank (Page et al., 1999), hub-authority scores (Kleinberg, 1999), in-degree and out-degree, betweenness (Freeman, 1977), closeness (Bavelas, 1948), and eigenvalue centrality (Bonacich, 1972).
Each centrality measure captures distinct aspects of student engagement in the network. PageRank evaluates student influence by considering both the number and importance of incoming connections, assigning higher scores to students who receive interactions from influential peers. Hub scores identify students who initiate connections with multiple peers, while authority scores highlight students who serve as primary information sources. In-degree and out-degree centrality measure the weighted volume of received and initiated interactions respectively, identifying key participants in collaborative discourse. Eigenvalue centrality assesses each student’s influence by considering the influence levels of their connected peers.
@@
  Angelina 0.1033 57 61 32 0.3695 0.2245 0.0378 0.0654
Armando 0.0414 46 46 34 0.4829 0.2242 0.1974 0.0198
Daniel 0.0286 48 37 27 0.4728 0.1811 0.0216 0.6029
Diego 0.0451 75 47 115 0.5616 0.5980 0.0147 0.7355
Esther 0.0479 92 105 76 0.5176 0.8422 0.0556 0.6152
Gio 0.0515 56 139 145 0.5454 0.2639 1.0000 0.0171
Isabel 0.0330 42 40 29 0.4775 0.2533 0.2659 0.0570
Jessica 0.0522 71 26 57 0.4916 0.7276 0.0133 0.5414
Julian 0.0530 87 51 52 0.5046 0.3090 0.0385 0.9111
Keila 0.0229 35 22 0 0.3728 0.1329 0.0080 0.0877
Kevin 0.0484 89 55 71 0.4571 0.2615 0.0045 1.0000
Kimberly 0.0481 51 93 51 0.4613 0.2644 0.1562 0.0420
Marco 0.0518 49 73 18 0.3276 0.1015 0.0067 0.2131
Maria 0.0607 41 53 61 0.4259 0.2476 0.1405 0.0570
Melissa 0.0184 25 46 0 0.3872 0.2847 0.1741 0.0217
Mya 0.0665 92 135 142 0.5677 1.0000 0.6806 0.0618
Natalie 0.0343 60 37 4 0.4005 0.4354 0.0421 0.2712
Ruben 0.0481 59 74 43 0.4086 0.1663 0.5060 0.0051
Samantha 0.0965 77 38 5 0.3629 0.2925 0.0247 0.0741
William 0.0485 62 36 7 0.3242 0.1290 0.1091 0.0133
Table 6: Centrality measures for the EXP (Explain Own Idea) network.
@@
  Angelina 0.0307 48 48 36 0.2968 0.0045 0.0006 0.0030
Armando 0.0431 96 50 9 0.6094 0.4061 0.0796 0.0262
Daniel 0.0358 36 33 0 0.4462 0.1607 0.0215 0.0174
Diego 0.0556 96 128 135 0.6766 0.5410 0.0292 0.4539
Esther 0.0964 192 75 110 0.6310 1.0000 0.0162 1.0000
Gio 0.0744 134 75 125 0.5837 0.3982 0.0477 0.0443
Isabel 0.0377 73 38 17 0.5011 0.2453 0.0999 0.0430
Jessica 0.0278 59 80 90 0.5569 0.1406 0.0098 0.1212
Julian 0.0600 71 152 57 0.5309 0.2287 0.0521 0.0290
Keila 0.0228 33 40 0 0.4228 0.1101 0.0055 0.0623
Kevin 0.0615 75 86 38 0.4894 0.2484 0.0398 0.0203
Kimberly 0.0537 73 97 63 0.5287 0.2807 0.2482 0.0173
Marco 0.0707 104 54 10 0.4116 0.2906 0.0092 0.0144
Maria 0.0275 51 77 68 0.5223 0.0788 0.0449 0.0060
Melissa 0.0481 59 58 11 0.5133 0.3470 0.3616 0.0274
Mya 0.0646 119 175 167 0.6586 0.6136 1.0000 0.0229
Natalie 0.0329 47 67 0 0.4311 0.1986 0.0909 0.1434
Ruben 0.0629 82 76 29 0.4609 0.2486 0.0209 0.0447
Samantha 0.0276 37 47 0 0.2647 0.0019 0.0044 0.0002
William 0.0662 71 100 20 0.4313 0.2460 0.0153 0.0144
@@
Table 7: Centrality measures for the EOI (Engage Others’ Idea) network.
Tables 6 and 7 illustrate how different centrality measures reveal distinct patterns of student engagement in the interaction networks. For example, in the explanation network, Mya demonstrates high PageRank centrality, indicating her frequent interactions with influential peers. Esther’s elevated in-degree and out-degree scores reflect her active participation as both recipient and initiator of mathematical explanations. Notably, Gio’s high betweenness centrality suggests a crucial role in facilitating information flow between different student groups, effectively bridging distinct classroom communities.
Analysis of Table 7 reveals distinct interaction patterns within the Engagement network, with several students demonstrating notable centrality profiles. Esther’s consistently high scores across centrality metrics indicate her comprehensive engagement in mathematical discourse, both influencing and responding to peers’ ideas. Gio maintains his bridging role while actively participating in peer discussions, while Mya demonstrates strong bilateral engagement through both contributing to and building upon others’ mathematical reasoning. Additionally, Armando, Diego, and Julian’s elevated eigenvector centrality and hub scores suggest their significant influence within localized peer groups, indicating the presence of distinct collaborative clusters within the classroom.
Our analysis shows that high network centrality does not necessarily correlate with improved mathematical performance. Mya’s case illustrates this complexity: despite her prominent position in both interaction networks, her mathematical achievement did not improve, possibly due to her frequent interactions with lower-performing peers. This suggests that the relationship between network centrality and academic performance is not straightforward and the specific composition of local interactions - such as which students interact with whom - appears to play a more significant role. The influence of classroom interactions on learning outcomes appears to depend more on the composition of immediate peer groups than on students’ overall network prominence. This observation underscores the importance of examining both the quality and context of student interactions within local network structures, rather than relying exclusively on global centrality metrics.
@@ 3.2 Network Mediation Model
Following the descriptive analysis of classroom interaction networks through visualization and centrality metrics, we now proceed to a more rigorous examination of how these network structures influence student outcomes. While network centrality measures offer valuable insights into the roles and prominence of individual students within the interaction network, they do not directly reveal the mechanisms through which these interaction patterns impact learning. To understand these mechanisms, we employ network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022). Network mediation analysis provides a framework for addressing this challenge by treating the network structure as a mediating variable. A key challenge in implementing network mediation analysis is representing the complex network structure in a way that can be used as a mediator. To address this, we adopt the approach of latent space model (Hoff et al., 2002; Hoff, 2005; Sewell and Chen, 2016, LSM)). LSM provide a framework to represent each student’s position within the network in a lower-dimensional latent space, capturing their underlying social dynamics wherein inter-student distances reflect the probability and intensity of their interactions.
@@
Our study examines how classroom interaction networks mediate the influence of gender on mathematics achievement. Prior research has consistently investigated gender differences in mathematics performance, revealing complex patterns and influences (Lindberg et al., 2010; Else-Quest et al., 2010; Ghasemi and Burley, 2019), while Lindberg et al. (2010) demonstrate that overall gender gaps in mathematics performance are relatively small. Ghasemi and Burley (2019) identifies significant cultural and societal influences on mathematics-related preferences and self-confidence across genders. Else-Quest et al. (2010) emphasizes the role of societal context in gender disparities, noting smaller performance differences in countries with greater gender equality. These findings collectively suggest that gender differences in mathematics achievement stem from social rather than purely biological factors, underscoring the importance of analyzing classroom social networks and their mediating effects on academic performance.
@@
Based on this perspective, we hypothesize that the gender gap in mathematics performance ($X$ on $Y$) can be differentiated based on how individual students interact with others during mathematics discussion classes. Specifically, we apply network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022), using students’ interaction structures as a mediator ($M$) for gender ($X$) - math performance ($Y$) relationships. By incorporating the estimated latent positions ($z_{i}$, $w_{j}$) as mediators, our goal is to investigate the indirect effects of gender on mathematics performance through the network structure.
Pre-test Scores as a Confounder
In the current context, pre-test scores could serve as a useful confounder for the intended mediation analysis. However, the pre-test (California Standardized test; CST) and post-test (a test developed by Webb et al. (2023)) do not measure the same domains, while the assessment methods and criteria are also different between the two tests. Thus, the two test scores may not be entirely comparable (Liu et al., 2017).
Figure 4: Scatter plot showing the relationship between pre-test (CST) and post-test scores.
Further, there is an additional technical challenge: the pre-test scores do not seem to have a linear relationship with the post-test scores. As shown in Figure 4, the relationship is not clearly linear around the Basic (300-350) and Proficient (350-400) ranges of the pre-test. This means that higher scores in the pre-test may not predict higher post-test scores consistently, suggesting pre-test scores may not be a reliable predictor of the post-test scores.
To account for students’ prior mathematical proficiency in our analysis, we constructed a binary proxy variable based on California Standards Test (CST) performance. Students who achieved scores of 350 or above — the state-designated threshold for proficiency [^2] https://www.starsamplequestions.org/about.html — were categorized as high-performing. This dichotomization aligns with state educational standards and provides a meaningful distinction in student achievement levels. We incorporate this proxy variable as a confounder in our mediation analysis to more precisely evaluate how network structures mediate gender-based differences in mathematics performance.
Figure 5: Network mediation model path diagram. This model analyzes the effect of gender on math achievement mediated by network interactions. Latent positions of sender and receiver networks are used as mediators.
@@
We additionally consider predictor-mediator interaction effects to account for potential heterogeneity in mediator effects across varying predictor levels (VanderWeele and Vansteelandt, 2009; Valeri and VanderWeele, 2013). In this case, the indirect effects, mediated through the latent network structure, vary with gender ($x$). By modeling this interaction, we account for the possibility that the indirect effect of gender on math performance is mediated differently depending on the students’ positions within the network. The resulting network mediation model can be expressed as:
@@
where $x$ is the binary indicator for gender, $m_{k}$ represents the latent positions of students, and $c$ is the confounder indicating proficiency based on the CST score threshold. Specifically, the latent variables $m_{k}$ represent the coordinates of the latent positions for the network nodes: $m_{1}$ and $m_{2}$ correspond to sender position coordinates (x and y axes, respectively), while $m_{3}$ and $m_{4}$ represent receiver position coordinates (x and y axes, respectively). The model coefficients capture distinct aspects of the gender-performance relationship: $a_{1k}$ quantifies gender’s influence on network positioning, $b_{2k}$ measures how network positions affect mathematics performance, and $b_{3k}$ captures the predictor-mediator interaction between gender and network position, revealing how network-mediated effects vary by gender. Additionally, $a_{2k}$ and $b_{4}$ account for the influence of prior mathematical proficiency on network positions and post-test performance, respectively. $a_{0k}$ and $b_{0}$ are intercept terms for the network mediation model.
This model enables the decomposition of the total effect (TE) into two distinct components: the natural direct effect (NDE) and the natural indirect effect (NIE). The NDE represents the component of the effect of gender ($X$) on math performance ($Y$) that is independent of the mediators, while the NIE captures the effect transmitted through the mediating network structure. These components constitute the total effect (TE), providing a comprehensive measure of the impact of gender on math performance. The natural direct effect, the natural indirect effect, and the total effect, are given by:
@@
Our analysis examines how distinct types of classroom interactions mediate gender-based differences in mathematics performance by investigating two separate networks: the EXP network and the EOI network. For both networks, we incorporate students’ prior mathematics achievement as a confounder to account for pre-existing academic differences. Through the application of our network mediation model to these distinct interaction networks, we aim to investigate whether the mediating effects of classroom network structures vary systematically based on interaction type.

### v2

@@ 2.4.3 Uncertainty Quantification
The resulting entropy values operate within a defined range from 0, indicating perfect consensus among all models, to $\log_{2}(n)$, representing maximum uncertainty where models are equally divided across all possible labels. Utterances exhibiting high entropy values signal substantial inter-model disagreement and are systematically flagged as priority candidates for expert human annotation. This entropy-driven selection strategy enables efficient allocation of limited annotation resources while maintaining rigorous quality standards, effectively reducing the computational and human burden associated with comprehensive manual review processes.
Figure 2: Distribution of Shannon entropy across seven LLMs for inter-model uncertainty quantification
Figure 2 illustrates the distribution of Shannon entropy values derived from classification results across our seven LLMs. The analysis of 4,233 total utterances reveals that approximately 57 percent (2,433 utterances) achieved complete consensus among all models, demonstrating substantial inter-model agreement that validates the reliability of our classification framework. The calculated 95th percentile entropy threshold of 1.38 represents classification scenarios where four of the seven models converge on identical label assignments, indicating that the vast majority of utterances generate strong consensus patterns across the ensemble. Conversely, the top 5 percent of utterances exhibiting the highest entropy values reveal significant inter-model disagreement, creating natural breakpoints where either appropriate post-processing techniques or expert human annotation becomes essential for resolving classification ambiguity. In this study, we employed majority voting among commercial LLM APIs as a post-processing approach, given their higher consistency compared to other models.
Intra-Model Uncertainty Quantification
To evaluate model confidence with greater precision, we implemented an intra-model uncertainty assessment with two distinct approaches. The first approach employed sequence-level probability analysis, which measures the probability of complete classification labels by computing normalized logit scores across all possible label sequences (e.g., “Explain Own Idea,” “Engage with Others’ Ideas” with varying engagement levels, or “Uncorrelated”) and then quantifies the overall confidence in the final classification decision. The second approach utilized token-level probability analysis, which examines the probability distribution at each token generation step throughout the classification process, effectively replicates the internal decision-making mechanism of the LLM and provides detailed insights into specific decision points where uncertainty emerges.
We applied Shannon entropy in Equation 2 as our primary uncertainty quantification metric for both analytical approaches. These complementary analytical approaches illuminate the distinct dimensions of model uncertainty, with sequence-level entropy that measures holistic linguistic confidence in complete label generation, and token-level entropy that identifies specific decision points where models demonstrate hesitation. Comprehensive sequence-level and token-level analysis results, model comparisons, and detailed methodological specifications are available in Section LABEL:supp-sec:uncertainty of the Supplementary Material.
2.4.4 Summary
Our comprehensive validation framework demonstrates the robustness and reliability of the LLM-based utterance classification approach through multiple analytical perspectives. Inter-model consistency analysis revealed substantial agreement levels across the ensemble, with commercial API models achieving the highest consensus ($\kappa_{Fleiss}$ = 0.7354) among all evaluated LLMs. Inter-model uncertainty quantification using Shannon entropy demonstrated that 57% of utterances achieved complete consensus across all models, while the top 5% with high uncertainty were effectively addressed through majority votes among commercial models. Intra-model uncertainty analysis provided complementary insights into model confidence patterns, revealing distinct uncertainty profiles across different model architectures.
Comparative analysis with human annotations yielded particularly significant insights regarding classification consistency patterns. LLM ensembles achieved markedly higher inter-annotator agreement ($\kappa_{Fleiss}$ = 0.5070-0.5933) compared to human annotators ($\kappa_{Fleiss}$ = 0.3509), highlighting the substantial variability inherent in human interpretation of dialogue data. This finding demonstrates the considerable value of automated classification approaches in addressing the inherent subjectivity challenges that characterize dialogue annotation tasks, where human judgments frequently vary due to individual interpretative differences and contextual perspectives, thus supporting the rationale for LLM-based classification in reproducible interaction analysis. These validation results collectively establish strong confidence in our classification framework’s ability to generate reliable and methodologically sound labels for subsequent network analysis applications.
@@ 3.1 Network Construction from Classified Utterances
We formalized student interactions using an adjacency matrix, where the matrix elements indicate the frequency of specific interaction types between two students during the observation period. We then visualized the constructed network structure using nodes to represent individual students and directed weighted edges between nodes to represent the nature and frequency of the interactions between the two students. For instance, when a student offers multiple explanations to a peer, a directed edge is formed from the explaining student (source) to the listening student (target), with the edge weight reflecting the total number of explanatory interactions. Engagement interactions similarly generate directed edges from the engaging student to the idea originator, with edge weights reflecting the interaction frequency. To capture the varying intensity of student engagement more precisely, we implemented a differential weighting scheme for these interactions. While explanatory interactions maintained equal weights, we applied multiplication factors to different levels of engagement: medium-level engagements received double weight, and high-level engagements received triple weight relative to low-level engagements. This weighted approach ensures that more substantive forms of participation exert proportionally greater influence on the structure of the network, better reflecting their educational importance.
@@ 3.1.1 Network Graphs
Figure 3 illustrates these network transformations, where each node represents a student, and the edges represent their interactions. The node colors indicate students’ standardized mathematics performance, utilizing a spectrum from blue (lower scores) to red (higher scores). The edge thickness corresponds to the interaction frequency between pairs of students.
@@
We then examined network centrality, using metrics designed for directed weighted networks, accounting for both directional flow and interaction intensity. These include PageRank (Page et al., 1999), hub-authority scores (Kleinberg, 1999), in-degree and out-degree, betweenness (Freeman, 1977), closeness (Bavelas, 1948), and eigenvector centrality (Bonacich, 1972).
Each centrality measure captures distinct aspects of student engagement in the network. PageRank evaluates student influence by considering both the number and importance of incoming connections, assigning higher scores to students who receive interactions from influential peers. Hub scores identify students who initiate connections with multiple peers, while authority scores highlight students who serve as primary information sources. In-degree and out-degree centrality measure the weighted volume of received and initiated interactions respectively, identifying key participants in collaborative discourse. Eigenvector centrality assesses each student’s influence by considering the influence levels of their connected peers.
@@
Angelina 0.0857 28 31 11 0.1725 0.2081 0.3438 0.1159
Armando 0.0223 8 8 0 0.1283 0.0694 0.0410 0.0258
Daniel 0.0229 10 8 0 0.1295 0.0129 0.0077 0.0199
Diego 0.0448 23 23 69 0.2064 0.4074 0.1418 0.4323
Esther 0.0730 54 38 46 0.2377 0.7694 0.3451 1.0000
Gio 0.0474 18 28 79 0.1828 0.0434 0.0718 0.0218
Isabel 0.0388 24 28 109 0.2384 0.0865 0.2362 0.2082
Jessica 0.0669 35 16 27 0.2047 0.6714 0.0684 0.7269
Julian 0.0513 28 28 115 0.2130 0.0430 0.0516 0.1074
Keila 0.0173 12 18 0 0.1916 0.0571 0.0765 0.1979
Kevin 0.0307 18 9 10 0.1327 0.0092 0.0030 0.0169
Kimberly 0.0517 38 54 123 0.2369 0.1693 0.5539 0.1953
Marco 0.0425 19 28 27 0.1569 0.0072 0.0134 0.0122
Maria 0.0548 15 23 7 0.1710 0.2063 0.2442 0.0445
Melissa 0.0198 10 15 0 0.1560 0.1180 0.2284 0.1004
Mya 0.1094 57 63 117 0.2340 1.0000 1.0000 0.4497
Natalie 0.0478 30 36 57 0.1870 0.3322 0.3351 0.4367
Ruben 0.0252 10 18 4 0.1243 0.0022 0.0059 0.0031
Samantha 0.0928 47 25 36 0.1845 0.2919 0.0892 0.5415
William 0.0548 34 21 55 0.1752 0.0200 0.0308 0.0284
Table 8: Centrality measures for the EXP (Explain Own Idea) network. The top three values in each column are highlighted in bold.
@@
Angelina 0.0354 42 38 36 0.1892 0.0020 0.0001 0.0054
Armando 0.0456 53 26 41 0.3819 0.1221 0.0618 0.0190
Daniel 0.0450 27 23 35 0.3289 0.1590 0.0026 0.0290
Diego 0.0412 68 68 126 0.4468 0.2104 0.0075 0.9682
Esther 0.0508 80 26 31 0.3907 0.1904 0.0063 1.0000
Gio 0.0878 82 63 148 0.4118 0.7956 0.0649 0.0100
Isabel 0.0554 48 22 17 0.3347 0.6473 0.0433 0.0228
Jessica 0.0225 35 32 69 0.3278 0.0437 0.0010 0.3052
Julian 0.1053 81 92 93 0.3811 1.0000 0.0109 0.0650
Keila 0.0400 38 35 0 0.2372 0.0445 0.0018 0.0723
Kevin 0.0454 37 44 30 0.2873 0.2871 0.0014 0.0174
Kimberly 0.0641 54 63 38 0.2682 0.0651 0.1144 0.0060
Marco 0.0585 69 44 5 0.2965 0.6780 0.0012 0.0251
Maria 0.0261 27 41 70 0.3044 0.0157 0.0497 0.0014
Melissa 0.0390 28 39 1 0.2515 0.0423 0.1496 0.0079
Mya 0.0507 54 139 167 0.4264 0.1348 1.0000 0.0064
Natalie 0.0472 51 41 47 0.2745 0.0613 0.0193 0.2643
Ruben 0.0462 41 42 3 0.2588 0.2731 0.0074 0.0010
Samantha 0.0313 30 39 0 0.1733 0.0010 0.0018 0.0001
William 0.0623 51 79 34 0.3228 0.6860 0.0263 0.0040
@@
Table 9: Centrality measures for the EOI (Engage Others’ Idea) network. The top three values in each column are highlighted in bold.
Tables 8 and 9 demonstrate how multiple centrality measures provide complementary insights into different patterns of student engagement within classroom interaction networks. The analysis reveals the differentiated roles that individual students take in facilitating mathematical discourse. Within the explanation network, Mya emerges as the most influential contributor, achieving the highest scores across the PageRank, out-degree, eigenvector, and hub centrality measures. On the other hand, Esther occupies a distinct but equally important position, ranking highest in in-degree and authority centrality scores. This pattern suggests that she serves as a focal point for peer explanations, with many students directing their mathematical reasoning toward her for consideration or validation. Kimberly demonstrates yet another critical network function, exhibiting the highest betweenness centrality score among all participants. This metric highlights her essential role as a connector who bridges other separate groups of students, thereby facilitating information flow and maintaining network cohesion throughout the classroom community.
Table 9 shows distinct interaction patterns in the engagement network that highlight the diverse roles of students in peer-to-peer mathematical discourse. Mya again emerges as the primary engagement initiator, achieving the highest scores in out-degree, betweenness, and hub centrality measures. These convergent metrics demonstrate her exceptional capacity for outgoing engagement and her pivotal position in the dissemination of ideas throughout the classroom network. Gio complements this pattern by demonstrating the highest PageRank score alongside top-tier performance in betweenness and degree centrality measures, establishing his critical function as a connector who bridges distinct subgroups within the classroom community. Esther ranks highest in in-degree and authority centrality while exhibiting comparatively lower out-degree and eigenvector scores. This profile indicates that she serves as a focal point for peer engagement, frequently receiving directed interactions from classmates despite maintaining a more selective approach to outward influence. Julian demonstrates influence through strategic positioning, achieving the top eigenvector centrality score alongside high-degree measures that suggest that his impact operates primarily through well-connected peer relationships rather than direct engagement volume. Melissa and Kimberly both show elevated hub scores that reflect their active participation in initiating discussions and amplifying peer contributions throughout the network.
Our exploratory network analysis suggests that high network centrality does not necessarily translate into improved mathematical performance, indicating that prominent positioning within classroom interaction networks alone may be insufficient to guarantee learning outcomes. Mya’s case exemplifies this complex dynamic. Despite maintaining prominent positions across explanation and engagement networks, her mathematical achievement remained stagnant throughout the study period. This counterintuitive result may be attributable to the composition of her peer interactions, particularly her frequent engagement with lower-performing students who may not have provided the mathematical scaffolding necessary for academic advancement. These findings demonstrate that the relationship between network position and academic success operates through more nuanced mechanisms than simple connectivity metrics suggest. The quality and academic caliber of immediate peer connections appear to exert greater influence on learning outcomes than overall network prominence or global centrality measures. This discovery underscores the critical importance of examining local network structures and interaction patterns rather than relying exclusively on aggregate centrality indicators. It is important to stress that the above interpretations, though, are exploratory solely based on network centrality measures, and do not necessarily imply causal relationships.
@@ 3.2 Network Mediation Model
Following the descriptive analysis of classroom interaction networks through visualization and centrality metrics, we now proceed to a more rigorous examination of how these network structures influence student outcomes. While network centrality measures offer valuable insight into the roles and prominence of individual students within the interaction network, they do not directly reveal the mechanisms through which these interaction patterns impact learning. To understand these mechanisms, we employ network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022). Network mediation analysis provides a framework for addressing this challenge by treating the network structure as a mediating variable. A key challenge in implementing network mediation analysis is representing the complex network structure in a way that can be used as a mediator. To address this, we adopt the approach of latent space model (Hoff et al., 2002; Hoff, 2005; Sewell and Chen, 2016, LSM)). LSM provide a framework to represent each student’s position within the network in a lower-dimensional latent space, capturing their underlying social dynamics wherein inter-student distances reflect the probability and intensity of their interactions.
@@
Our study examines how classroom interaction networks mediate the influence of gender on mathematics achievement. Previous research has consistently investigated gender differences in mathematics performance, revealing complex patterns and influences (Lindberg et al., 2010; Else-Quest et al., 2010; Ghasemi and Burley, 2019), while Lindberg et al. (2010) demonstrate that overall gender gaps in mathematics performance are relatively small. Ghasemi and Burley (2019) identifies significant cultural and societal influences on mathematics-related preferences and self-confidence across genders. Else-Quest et al. (2010) emphasizes the role of the social context in gender disparities, noting smaller performance differences in countries with greater gender equality. These findings collectively suggest that gender differences in mathematics achievement stem from social rather than purely biological factors, underscoring the importance of analyzing classroom social networks and their mediating effects on academic performance.
@@
Based on this perspective, we hypothesize that the gender gap in mathematics performance ($X$ on $Y$) can be differentiated based on how individual students interact with others during mathematics discussion classes. Specifically, we apply network mediation analysis (Liu et al., 2021; Che et al., 2021; Di Maria et al., 2022), using students’ interaction structures as mediators ($M$) for gender ($X$) - math performance ($Y$) relationships. By incorporating estimated latent positions ($z_{i}$, $w_{j}$) as mediators, our goal is to investigate the indirect effects of gender on mathematics performance through the network structure.
To address potential confounding effects, we incorporated students’ prior mathematical proficiency as a confounding variable in our mediation analysis. We use California Standards Test (CST) performance, with students scoring 350 or above the state-designated proficiency threshold classified as high-performing and those below this threshold categorized as requiring additional support. This approach addresses non-linear relationships observed between pre-test and post-test performance measures while systematically controlling for baseline achievement differences that could otherwise confound the relationship between gender, network position, and learning outcomes. Detailed justification for our choice of the confounding variable is provided in Section LABEL:supp-sec:confounding of the Supplementary Material.
@@
We also consider the effects of the predictor-mediator interaction to account for the potential heterogeneity in mediator effects at different predictor levels (VanderWeele and Vansteelandt, 2009; Valeri and VanderWeele, 2013). In this case, indirect effects, mediated through the latent network structure, vary by gender ($x$). By modeling this interaction, we account for the possibility that the indirect effect of gender on math performance is mediated differently depending on the students’ positions within the network. The resulting network mediation model can be expressed as follows:
@@
where $x$ is the binary indicator for gender, $m_{k}$ represents the latent positions of the students, and $c$ is the confounder indicating proficiency based on the CST score threshold. Specifically, the latent variables $m_{k}$ represent the coordinates of the latent positions for the network nodes: $m_{1}$ and $m_{2}$ correspond to the coordinates of the sender position (x and y axes, respectively), while $m_{3}$ and $m_{4}$ represent the coordinates of the receiver position (x and y axes, respectively). The model coefficients capture distinct aspects of the gender-performance relationship: $a_{1k}$ quantifies gender’s influence on network positioning, $b_{2k}$ measures how network positions affect mathematics performance, and $b_{3k}$ captures the predictor-mediator interaction between gender and network position, revealing how network-mediated effects vary by gender. Additionally, $a_{2k}$ and $b_{4}$ account for the influence of prior mathematical proficiency on network positions and post-test performance, respectively. $a_{0k}$ and $b_{0}$ are intercept terms for the network mediation model.
This model enables the decomposition of the total effect (TE) into two distinct components: the natural direct effect (NDE) and the natural indirect effect (NIE). The NDE represents the component of the effect of gender ($X$) on math performance ($Y$) that is independent of the mediators, while the NIE captures the effect transmitted through the mediating network structure. These components constitute the total effect (TE), providing a comprehensive measure of the impact of gender on math performance. The natural direct effect, the natural indirect effect, and the total effect are given by:
@@
Our analysis examines how distinct types of classroom interactions mediate gender-based differences in mathematics performance by investigating two separate networks: the EXP network and the EOI network. For both networks, we incorporate students’ prior mathematics achievement as a confounder to account for preexisting academic differences. Through the application of our network mediation model to these distinct interaction networks, we aim to investigate whether the mediating effects of classroom network structures vary systematically based on interaction type.